# Searching Meta Reasoning Skeleton to Guide LLM Reasoning

### v1

@@ A.3 Overall Search Algorithm
We show the overall search algorithm in Algorithm 2. We set implement $N$ as 8, $M$ as 16 and learning rate $\eta$ to $5\times 10^{-4}$ during search for both tasks. We refer to previous works (Zhuge et al., 2024; Zhang et al., 2025a; Xie et al., 2025; Hu et al., 2025; Cheng et al., 2020), implement techniques such as gradient clipping, to improve the stability and convergence rate of search algorithm. See our code for implementation details. We implement a rule-based $r$ by exactly matching final answer $\hat{a}=\text{LLM}(q,\alpha_{q})$ given by LLM with ground-truth $a$ from dataset. Specifically,
@@ C.1 Baseline Implementation
  • MRP. MRP dose not have open-source code, but provides prompt in original paper. We follow the paper to implement MRP.
  • Meta-Reasoner. Meta-Reasoner dose not have open-source code, but provides prompt, pseudo code and detailed description in original paper. We follow the paper to implement Meta-Reasoner.
@@
  • RS. Referring to previous works (Bergstra & Bengio, 2012; Liu et al., 2019), we sample 48 architectures from search space randomly. Then we validate these architectures on training set to select the one with highest accuracy. With the selected architecture, we report its accuracy on test set.
  • QI. Referring to previous works (Liu et al., 2019; Zhuge et al., 2024), we do not use an MLP which takes reasoning context as input and output meta strategy distribution, but model the strategy distribution of each edge in search space without condition. We optimize the distribution with the same estimation of policy gradient with REINFORCE as in Equation 4. For all queries in test set, we sample only one skeleton to process all of them.
  • CA. Referring to previous works (Cheng et al., 2020; Zhang et al., 2025a), we use an MLP which takes semantic embedding of queries and meta reasoning strategies existing in skeleton as input to sample strategy for edges, rather than based on base reasoning context. For each query in test set, we sample a complete skeleton before inference and then reason for the query guided by the complete skeleton.
@@ C.2 Datasets Details
For training set, we use MATH  [^5] https://github.com/hendrycks/math training split composed of 5053 query-answer pairs and MMLU-Pro  [^6] https://github.com/TIGER-AI-Lab/MMLU-Pro training split composed of 70 query-answer pairs. For testing set, we use GSM8K  [^7] https://github.com/openai/grade-school-math, MATH-500  [^8] https://huggingface.co/datasets/HuggingFaceH4/MATH-500, AMC  [^9] https://huggingface.co/datasets/AI-MO/aimo-validation-amc, Olympiad  [^10] https://github.com/OpenBMB/OlympiadBench and four subset (Science, Humanities, Social and Other) of MMLU-Pro. We summarize the statistics of dataset in Tabel 3.

### v2

@@ A.3 Overall Search Algorithm
We show the overall search algorithm in Algorithm 2. We set implement $N$ as 8, $M$ as 16 and learning rate $\eta$ to $5\times 10^{-4}$ during search for both tasks. We refer to previous works [zhuge2024gptswarm, zhang2025multiagent, logicrl, reinforce++, cheng2020instanas], implement techniques such as gradient clipping, to improve the stability and convergence rate of search algorithm. See our code for implementation details. We implement a rule-based $r$ by exactly matching final answer $\hat{a}=\text{LLM}(q,\alpha_{q})$ given by LLM with ground-truth $a$ from dataset. Specifically,
@@ C.1 Baseline Implementation
  • MRP. MRP does not have open-source code, but provides prompt in original paper. We follow the paper to implement MRP.
  • Meta-Reasoner. Meta-Reasoner does not have open-source code, but provides prompt, pseudo code and detailed description in original paper. We follow the paper to implement Meta-Reasoner.
@@
  • RS. Referring to previous works [randomsearch, darts], we sample 48 architectures from search space randomly. Then we validate these architectures on training set to select the one with highest accuracy. With the selected architecture, we report its accuracy on test set.
  • QI. Referring to previous works [darts, zhuge2024gptswarm], we do not use an MLP which takes reasoning context as input and output meta strategy distribution, but model the strategy distribution of each edge in search space without condition. We optimize the distribution with the same estimation of policy gradient with REINFORCE as in Equation 4. For all queries in test set, we sample only one skeleton to process all of them.
  • CA. Referring to previous works [cheng2020instanas, zhang2025multiagent], we use an MLP which takes semantic embedding of queries and meta reasoning strategies existing in skeleton as input to sample strategy for edges, rather than based on base reasoning context. For each query in test set, we sample a complete skeleton before inference and then reason for the query guided by the complete skeleton.
@@ C.2 Datasets Details
For training set, we use MATH  [^5] https://github.com/hendrycks/math training split composed of 5053 query-answer pairs and MMLU-Pro  [^6] https://github.com/TIGER-AI-Lab/MMLU-Pro training split composed of 70 query-answer pairs. For testing set, we use GSM8K  [^7] https://github.com/openai/grade-school-math, MATH-500  [^8] https://huggingface.co/datasets/HuggingFaceH4/MATH-500, AMC  [^9] https://huggingface.co/datasets/AI-MO/aimo-validation-amc, Olympiad  [^10] https://github.com/OpenBMB/OlympiadBench and four subset (Science, Humanities, Social and Other) of MMLU-Pro. We summarize the statistics of dataset in Tabel 4.
@@
C.3 Additional Results on Larger LLM
Besides Qwen2.5-3B-Inst and LLaMA3.2-3B-Inst used in Section 4.1 with token budget 1024, we conduct additional experiments to evaluate AutoMR based on larger long-cot LLM, including Qwen3-8B and Qwen3-14B (both enabling thinking mode, refer to Qwen3 homepage  [^11] https://qwenlm.github.io/blog/qwen3/ for more details about thinking mode) with larger token budget 8192. Additionally, we also evaluate on Pangu-7B with token budget 4096 on Ascend. We evaluate AutoMR and two meta reasoning methods that performs well in Table 1 on two datasets (MATH500 and Olympiad).
The results are showed in Table 5, where Base means the thinking mode of Qwen3 and Pangu7B with CoT. These results show that AutoMR still achieves performance gains when scaling to larger LLMs with larger token budget, compared with naive LLM thinking and previous meta reasoning methods, demonstrating the scaling potential of AutoMR.
Table 5: Additional results on Qwen3-8B, Qwen3-14B and Pangu-7B.
Method MATH-500 Olympiad
Pangu-7B Qwen3-8B Qwen3-14B Pangu-7B Qwen3-8B Qwen3-14B
Base (S) 75.4 80.8 84.2 48.5 47.0 49.0
Meta-Reasoner(S) - 81.6 85.6 - 49.9 50.4
rStar (T) - 83.4 84.8 - 48.9 50.9
AutoMR (G) 77.4↑2.0 86.6↑5.8 87.2↑3.0 49.2↑0.7 53.9↑6.9 54.5↑5.5